40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6809     Entry time: Tue Jun 12 23:18:18 2012
Author: yuta 
Type: Update 
Category: Green Locking 
Subject: I-Q signals for the beat 

[Mengyao, Yuta]

Yes!! We have I-Q signals for the beat!!

What we did:
  1. Aligned Y arm to the Y end green incident beam. The transmission to the PSL was about 195 uW.

  2. Aligned IR beam to the Y arm by adjusting PZTs and got the transmission, C1:LSC-TRY_OUT ~ 0.86.

  3. Aligned green optics on the PSL table to get the beat signal. The beat was found when;

  PSL laser temperature on display: 31.41 deg C
  C1:PSL-FSS_SLOWDC = 1.43
  Y end laser "T+": 34.049 deg C
  Y end laser "ADJ": 0
  Y end laser measured temperature: 34.14 deg C
  C1:GCY-SLOW_SERVO2_OFFSET = 29950
  Y end slow servo: off (was on)

  4. Connected the beat PD output to the beatbox.

  5. Kicked ETMY position to change the cavity length and while the ringdown, we run pynds to get data. We plotted C1:ALS-BEATY_FINE_I_ERR vs C1:ALS-BEATY_FINE_Q_ERR, and C1:ALS-BEATY_COARSE_I_ERR vs C1:ALS-BEATY_COARSE_Q_ERR (below). We got nice circle as expected.

FINEIQplot20120612.pngCOARSEIQplot20120612.png

Current setup:
  Only AA filers are put between the output of the beatbox and the ADC.

beatysetup20120612.png

ELOG V3.1.3-